### v6

@@ I Introduction
End-to-end learning has become a preferable approach in autonomous driving as manual configuration to integrate task-specific modules is no longer needed. This technique allows the model to share useful features directly from perception modules to controller modules. Moreover, the model can learn and receive extra supervision from a multi-task loss function that considers several performance criteria. All these benefits result in a better model performance even with a smaller model size due to its compactness [1][2]. To date, there have been a lot of works in the field of end-to-end autonomous driving, whether it is based on simulation [3], or offline real-world where the model predicts a set of driving records [4], or online real-world where the model is deployed for automated driving [5]. Besides dealing with diverse conditions such as driving on a sunny day or low-light evening, another challenge that remains in online real-world autonomous driving is that the model must deal with noise and inaccuracy of sensor measurement. This issue needs to be addressed as it affects model performance [6][7].
@@ III-A Proposed Model
       steering $=\mathbf{MLP}_{ST}$; throttle $=\mathbf{MLP}_{TH}$
@@
       steering $=\mathbf{PID}_{ST}$; throttle $=\mathbf{PID}_{TH}$
@@ III-B Model Improvement
Different from our previous work [1], the model is modified to improve its performance and deal with real-world implementation issues. First, as the input to the perception module, we consider a wider ROI of $H\times W=512\times 1024$ at the center of the RGBD image. Then, they are resized to $H\times W=256\times 512$ to reduce the computational load. With a wider input resolution, the model is expected to have a better scene understanding capability. Second, as the input to the controller module, we feed the left and right wheel’s angular speeds. This information is expected to be helpful, especially during turning as the angular speed will be different on each wheel. Third, two route points are given instead of one route point at a time. Relying only on one route point is very risky due to the possibility of mislocation caused by GNSS and IMU inaccuracies that affect the global-to-local coordinate transformation. If the route point is mislocated, the model will likely fail to predict waypoints and navigational controls correctly. Besides that, giving two route points will give the model a better intuition in deciding whether the vehicle should drive straight or turn depending on the location of the route points. Fourth, we also modify the control policy by allowing an agent to take the steering control completely over the other agent for better maneuverability.
@@ III-C Dataset
Conditions
Noon and evening
Total routes
12 (train-val) and 6 (test)
$\mathcal{N}$ samples*
10151 (train), 9679 (val), 18975 (test)
Devices
WHILL C2 vehicle (+rotary)
@@
Object classes
None, road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, terrain, sky, person, rider, car, truck, bus, train, motorcycle, bicycle
@@ III-E Evaluation and Scoring
  • The best performance is defined by the lowest total metric score as formulated with (10). $IoU_{SEG}$: IoU score of semantic segmentation. $MAE_{DE}$: mean absolute error of normalized depth estimation. $MAE_{WP}$: mean absolute error of waypoints prediction. $MAE_{ST}$: mean absolute error of steering estimation. $MAE_{TH}$: mean absolute error of throttle estimation. Latency: the time (in seconds) required by the model to process one set of observation data. To ensure the latency calculation is conducted fairly, we run each model independently on the same computer with NVIDIA RTX 3090.

### v7

@@ I Introduction
End-to-end learning has become a preferable approach in autonomous driving as manual configuration to integrate task-specific modules is no longer needed. This technique allows the model to share useful features directly from perception modules to controller modules. Moreover, the model can learn and receive extra supervision from a multi-task loss function that considers several performance criteria. All these benefits result in a better model performance even with a smaller model size due to its compactness [1][2]. To date, there have been a lot of works in the field of end-to-end autonomous driving, whether it is based on simulation [3], or offline real-world where the model predicts a set of driving records [4], or online real-world where the model is deployed for automated driving [5]. Besides dealing with diverse conditions such as driving on a sunny day or low-light evening, another challenge that remains in online real-world autonomous driving is that the model must deal with noise and inaccuracy of sensor measurement. This issue needs to be addressed as it affects model performance and its generalization capability [6][7].
@@ III-A Proposed Model
       steering $=\mathbf{MLP}_{ST}$
       throttle $=\mathbf{MLP}_{TH}$
@@
       steering $=\mathbf{PID}_{ST}$
       throttle $=\mathbf{PID}_{TH}$
@@ III-B Model Improvement
Different from our previous work [1], the model is modified to improve its performance and deal with real-world implementation issues. First, as the input to the perception module, we consider a wider ROI of $H\times W=512\times 1024$ at the center of the RGBD image. Then, they are resized to $H\times W=256\times 512$ to reduce the computational load. With a wider input resolution, the model is expected to have a better scene understanding capability. Second, as the input to the controller module, we feed the left and right wheel’s angular speeds. This information is expected to be helpful, especially during turning as the angular speed will be different on each wheel. Third, two route points are given instead of one route point at a time. Relying only on one route point is very risky due to the possibility of sensor inaccuracies that affect the global-to-local coordinate transformation. Although we have used RTK-GNSS and achieved 1cm positional accuracy, the coordinate transformation can be still disturbed by the inaccuracy of the IMU sensor. If the route point is mislocated, the model will likely fail to predict waypoints and navigational controls correctly. Besides that, giving two route points will give the model a better intuition in deciding whether the vehicle should drive straight or turn depending on the location of the route points. Fourth, we also modify the control policy by allowing an agent to take the steering control completely over the other agent for better maneuverability.
@@ III-C Dataset
Conditions Noon and evening
Total routes 12 (train-val)
6 (test)
$\mathcal{N}$ samples* 10151 (train) and 9679 (val)
18975 (test)
Devices WHILL C2 vehicle
@@
WHILL’s rotary encoder
Object classes None, road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, terrain, sky, person, rider, car, truck, bus, train, motorcycle, bicycle
@@ III-E Evaluation and Scoring
  • The best performance is defined by the lowest total metric score as formulated with (10). $IoU_{SEG}$: IoU score of semantic segmentation. $MAE_{DE}$: mean absolute error of normalized depth estimation. $MAE_{WP}$: mean absolute error of waypoints prediction. $MAE_{ST}$: mean absolute error of steering estimation. $MAE_{TH}$: mean absolute error of throttle estimation. Latency: the time (in seconds) required by the model to process one set of observation data. To ensure a fair latency calculation, we run each model independently on the same PC with NVIDIA RTX 3090.